OpenStudy (anonymous):

No idea how to solve this. Please help? What is the pressure in a 200.0 mL flask (at T = 27°C) after Ar from a 0.600 L container at 0.600 atm and 227°C is mixed with oxygen from a 0.400 L container at 1.20 atm and 127°C? Assume no reaction takes place between the argon and oxygen.

OpenStudy (somy):

well im guessing you would just add up pressures by each gas

OpenStudy (anonymous):

The answer is 2.88 atm but I dont know how to solve it

OpenStudy (somy):

wait a sec

OpenStudy (somy):

yay got it!

OpenStudy (somy):

alright so use PV= nRT to find moles of Ar & O n= PV / RT P unit atm V unit L R = 0.08206 T unit Kelvin u have all needed values for both Ar & O to calculate their moles

OpenStudy (somy):

so, once you get the moles add them up and you will get total moles

OpenStudy (somy):

now use PV= nRT again & this time u will find pressure P = nRT / V n= your total mole R = 0.08206 T = 27 + 273 = K V = 200 mL / 1000 = L so use the formula & get your final pressure in atm

OpenStudy (anonymous):

I need to find the pressure though, not the moles

OpenStudy (somy):

ofc, but you need to find moles first

OpenStudy (somy):

try working out whatever i told you you will get your answer

OpenStudy (anonymous):

Thank you!

OpenStudy (somy):

no problem :)